fontexport(1) teTeX fontexport(1) NAME fontexport−exportfontsfromtheteTeXdirectorystructure SYNOPSIS fontexport[−d]destinationfiles... DESCRIPTION fontexportexportsfontfilesfromtheteTeXdirectorytree. fontexportcancopewithfontnamesconformingthe8.3−namescheme,too,justgiveitthe −d−flag(das"dos").SoyoucanexportyourteTeXfontsforsomeoneelsetousethemwithhis orheremTeX. SEEALSO fontimport(1). AUTHOR ThomasEsser(te@informatik.uni−hannover.de) teTeX 11/94 1